OpenStudy (anonymous):

Show that there is no integer a such that a^2 − 3a − 19 is divisible by 289

OpenStudy (anonymous):

problemsolver where do u get all these problems ?

OpenStudy (anonymous):

maybe you could show that: \[a^2-3a-19 = 289k, k\in\mathbb Z\] is impossible? a proof by contradiction?

OpenStudy (anonymous):

ya contradiction

OpenStudy (anonymous):

i dont find prblems they find me

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

\[a^2-3a-19 = 289k \Rightarrow a^2-3a-(19+289k) = 0\] Quadratic formula: \[\frac{3\pm \sqrt{9+4(19+289k)}}{2}\] to show this has no integer solutions, i just want to show that the discriminant cant be a perfect square......right?

OpenStudy (anonymous):

don no wat u mean sry

OpenStudy (anonymous):

i want to show that the inside of that square root cant be a perfect square. i think thats what im trying to say lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

80 +1156k inside root

OpenStudy (anonymous):

so: \[9+4(19+289k) = 9+76+1156k = 85+1156k\] hmmm.....

OpenStudy (anonymous):

ah, a 17 factors out, kinda convenient see that 289 is 17^2 <.< so now we have: \[17(5+68k)\]

OpenStudy (anonymous):

good morning satellite!

OpenStudy (anonymous):

i think the idea is to see what happens when you square numbers of the form \[17j\] \[17j+1\] \[17j+2\] etc. but the only reason i think that is because i vaguely remember this, so perhaps this is cheating. good morning joemath!

OpenStudy (anonymous):

now , if the inside of the square root were a perfect square, then (5+68k) should also have a factor of 17 in it: \[5+68k \equiv 0 \mod 17\] but, 68 = 17*4, so we have: \[5+68k = 5+17*4k \equiv 5 \mod 17\] There is no way for a factor of 17 to come out of that expression. Therefor the inside of the square root will never be a perfect square, and there are no integer solutions blah blah contradiction blah.